Four pencils and two pens cost $0.74. Six pencils and five Pena cost $1.53. Find the cost of a pencil and a pen
Answer by mananth(12270)

Four pencils and two pens cost $0.74. Six pencils and five Pena cost $1.53. Find the cost of a pencil and a pen.
..
For easier caculation lets solve in cents
let pencil cost x cents
pen cost y cents
4x+2y=74.....................1
....
6x+5y=153 ...................2
..
multiply 1 by -6 & 2 by 4
-24x-12y=-444
24x+20y=612
add them
8y=168
/8
y= 21 cents, cost of pen
..
plug value of y in 1
4x+2*21= 74
4x+42=74
4x= 74-42
4x=32
/4
x= 8 cents ,cost of pencil
...
m.ananth@hotmail.ca